# revtex4-2Repair the float

### v1

@@ II.1 Setting and conventions
Figure 1: Cycle benchmarking circuit design [14] leveraged to learn individual Pauli eigenvalues to relative precision. Illustrated in this example is the simpler form that is sufficient for any Clifford hard cycle $\mathcal{H}$ where easy cycles are Pauli randomized cycles $P$  [15]. The easy cycle $B$ includes local rotations to prepare Pauli product states and define the measurement basis. The randomized Pauli cycle $A$ eliminates bias from measurement errors.
@@
We will also consider the restriction of elements in $\mathbb{P}_{n}$ to arbitrary subsets of qubits. The restriction of Pauli $P$ to the qubits indexed by elements of set $S$ is denoted $P_{S}$. For example, $[Z_{1}X_{2}Y_{4}]_{\{1\}}=Z_{1}$, $[Z_{1}X_{2}Y_{4}]_{\{2,3\}}=X_{2}I_{3}$, or $[Z_{1}X_{2}Y_{4}]_{\{4,1\}}=Y_{4}Z_{1}$. Note that the ordering of indices in the set is used here to typographically order the restricted Pauli. Similarly, we denote by $\mathbb{P}_{S}$ the subgroup of $\mathbb{P}_{n}$ that is supported on only the qubits of $S$. It will also be useful in this context, to define the map $\omega$ between arbitrary Pauli operators $P$ and $Q$, defined as
@@
where $p_{i}=\chi_{ii}$ is the classical probability of some Pauli operator $P_{i}$ occurring in the quantum channel. Evidently, the number of non-trivial parameters in $\mathcal{P}$ is drastically smaller than in $\mathcal{G}$; at most $14^{n}-$ 1. These smaller operators are often more practical to study.
@@
Channels of interest are generally hard cycles that are not the identity or idling operation. Typically we have some ideal channel $\mathcal{H}$ that is implemented imperfectly as $\tilde{\mathcal{H}}=\mathcal{E}\circ\mathcal{H}$ where $\mathcal{E}$ is the noise channel we are interested in quantifying. We are thus not always able to insert random Pauli operators either side of the noise channel $\mathcal{E}$, but we may still interleave operations either side of $\tilde{\mathcal{H}}$. We can effectively twirl $\mathcal{E}$ if we instead insert a Pauli operator $P_{i}^{c}=\mathcal{H}^{\dagger}(P_{i})$ before $\tilde{\mathcal{H}}$ for each random $P_{i}$ inserted after $\tilde{\mathcal{H}}$.
@@ II.3 Probability Marginalization
$\mu_{S}(P)=\sum_{\begin{subarray}{c}P_{i}\in\mathbb{P}_{n}\\ [P_{i}]_{S}=P\end{subarray}}p(P_{i}).$ (7)
Generically we refer to a marginal probability distribution over $n$ qubits as ‘$n$-qubit’ marginals.
@@
[…]
@@ II.4 Graphical model under Gibbs random field
The error correlations of single-qubit and entangling gates realized in the used trapped-ion experiment are assumed to be nearest-neighbor. We can abstract to a Gibbs random field connectivity graph as shown in Fig. 4 c, which we will later use to make predictions about logical performance. Each qubit is connected to its partner in the entangling gate and the corresponding neighbors.
The considerations above are tailored to our experiment and are not true in general. Other platforms may be dominated by different error sources and error correlations. As long as the error correlations are finite, analog GRF models can be derived and the calculations of logical performance adapted.
@@ II.5 Optimizing Experimental Parameters
In the same vein, we want to learn chosen experimental parameters in as few experimental runs as possible. One can extract up to $2^{n}$ different Pauli expectation values in parallel, but generally this requires entangled input states and measurements which contribute to SPAM. However, for hard cycles from the Clifford group certain finite sequence lengths are equivalent to performing the identity operation. Our hard cycle is a transversal application of cnots, so we specifically have $\mathcal{H}^{2}(\rho)=\rho$. Hence, by choosing $\mathcal{H}^{m}$ such that $m$ is even, we guarantee that easily-prepared Pauli product states will be mapped to Pauli product measurements, keeping the contribution of SPAM low. At the same time, the measurements resolve each qubit and we can infer multi-qubit Pauli measurement outcomes by taking the product of single-qubit Pauli measurement outcomes. This reduces the numnber of experiments further because we are able to extract all $2^{n}$ distinct Pauli expectation values from $n$-qubit product states.
Figure 3: Optimization of experimental resources. a The number of distinct experiments to run for a single, two, and seven parallel cnots to obtain all unique orbits is 4 (the eigenstates of elements from the set $A$, for 1-cnot orbits), $4\times 9=36$ (all combinations of $A$ and $B$, for 2-cnot orbits), and $36\times 3=108$ (all tilings $C$ of elements from $A$ and $B$, for all possible 2-cnot orbits over 7 cnots), respectively. b Tradeoff between number of randomizations per initial state and shots per randomization for an eigenvalue estimate of $\lambda_{ZX}$. The color scale depicts the standard deviation in estimates of $\lambda_{ZX}$ at the chosen experimental parameters, estimated by subsampling over an experimental data set. In total the data set included 300 randomizations with 650 shots each. In our platform changing the randomization is more costly than repeating a shot. Time in minutes to estimate a fixed initial state using three values for $m$ is indicated by red contours. Total shots required is obtainable by the product of the shots per randomization, total randomizations, the distinct sequence lengths $m$, and the distinct initial states.
@@
A naive estimate for the total number of configurations needed for 2-cnot marginals would be $4\times 4=16$ initial states. However, the set $(P_{1}P_{2})^{\circlearrowright\mathcal{H}}\neq P_{1}^{\circlearrowright\mathcal{H}}\times P_{2}^{\circlearrowright\mathcal{H}}$. This can be observed simply by noting the cardinality of the sets: $(P_{1}P_{2})^{\circlearrowright\mathcal{H}}=\{P_{1}P_{2},\mathcal{H}(P_{1}P_{2})\}$ has an orbital length of two but $P_{1}^{\circlearrowright\mathcal{H}}\times P_{2}^{\circlearrowright\mathcal{H}}$ is a set of 4 elements. For our choice of hard cycle, equality is satisfied whenever $P_{1}$ or $P_{2}$ is mapped to itself, up to a change in sign. Consequently we often need to prepare $P_{1}$ and $P_{1}P_{2}$, increasing the number of initial states from 16 to 36. When considering all possible 2-cnot marginalizations in the transversal seven cnot case, three distinct tilings, shown in Fig. 3 a, are needed such that 100 initial states are required in total ($3\times 36=108$, minus 8 redundancies) to learn $53764^{4}\times\binom{7}{2}=$ 5376 expectation values.
@@ II.6 Uncertainties and noise floors
Error bars shown in figures depicting marginal error rates are determined by bootstrapping the data over the time period of the experiment. Where calculated, the range is reported to one standard error unless stated otherwise.
@@ III Experimental setup
[…]
The native gate set used in the setup consists of arbitrary-angle single-qubit rotations $R(\theta,\phi)$ around any axis in the equatorial plane of the Bloch sphere (pulse area $\theta$ and angle $\phi$ relative to $+X$) implemented via laser pulses resonant with the optical qubit transition, single-qubits $Z$-rotations ‘virtually’ implemented in software, and maximally-entangling two-qubit gates $MS(\pm\pi/2)=\mathrm{exp}(\mp i\frac{\pi}{4}X\otimes X)$ implemented via the Mølmer-Sørensen interaction [41]. The $MS(\pm\pi/2)$ gate is equivalent to the cnot gate up to local rotations [42], compiled here as
@@
The $MS$ gates utilize the center-of-mass radial motional mode of the ion crystal. Using the global phononic bus of the motion to create entanglement means that we have to perform $MS$ gates serially even when two gates feature disjoint support. There are control-theoretic gate modulation techniques to circumvent this limitation, e.g. in Ref. [43], which are however not part of this work.
@@ IV.2 Transversal CNOT
Figure 6: Characterizations of a 14-qubit transversal cnot $\mathcal{H}=\prod_{i=0}^{6}\textsc{cnot}_{i,i+9}$ in a 16-ion register. a Heatmap display of 1-cnot marginals for each constituent cnot. The $ZX$ error terms are identified to be caused by a miscalibration of the $MS$ gate power. b A heatmap display of 1-cnot marginals for a separate characterization experiment, after the source of the $ZX$ signal was identified and eliminated. In a and b, marginals $\mu_{S}(ZI^{\circlearrowright\mathcal{H}})$ and $\mu_{S}(IZ^{\circlearrowright\mathcal{H}})$ associated with dephasing are larger than for a single cnot in isolation. The error characteristics for each of the individual cnots is similar (little change horizontally) with the exception of cnot0,9 for $\mu_{S}(IY^{\circlearrowright\mathcal{H}})$ and $\mu_{S}(IZ^{\circlearrowright\mathcal{H}})$ (highlighted in red). c Significant error terms for a 2-cnot marginal, taken at the same time as b, over cnot5,14 and cnot6,15. Marginal terms only supported on $(6,15)$ are shown first, then terms only supported on $(5,14)$, and then all other terms with error rates higher than the dark-gray region. d The same seven orbits from b as histograms, with the anomalous strengths of cnot0,9 again highlighted in red. Apart from the anomalous terms in cnot0,9, there is a much higher amount of dephasing on control and target qubits due to time spent idling. Since the predominant difference from single cnots in isolation appears to be idling, we compare the noise profile of individual cnots from the transversal cnot of d with a single cnot undergoing an equivalent period of idling. e Experimental sequence for labels ‘delay’ and ‘seven’ in f. The top circuit executes seven cnots successively and we study the marginal on support $(0,9)$ and $(3,12)$. The bottom circuit executes a single cnot but idles for an additional time equivalent to 6 cnots being applied in sequence. f and g are histograms of the marginals over cnot0,9 and cnot3,12, respectively, either as part of the full transversal cnot or as a single cnot with an equivalent idling time of the transversal cnot. Data in g was taken with machine performance deteriorated before ongoing larger maintenance. The idling errors of f resemble the transversal cnot errors, rather than additional dephasing.
@@
The overall wall-clock runtime for CER of a transversal cnot suite with 20 randomizations per initial states and 200 shots per randomization on seven supports takes about 40 hours after optimization; this time includes automatic calibration routines and other overhead of operation.
@@ IV.3 Predicting logical error rates and performance
Figure 7: Probabilities of different error scenarios reconstructed from transversal cnot CER data using the GRF factor graph of Fig. 4 c. Reconstructed non-zero error rates are grouped into uncorrectable and correctable errors (including errors stabilizer-equivalent to correctable, see Sec. II.4) assuming two logical qubits encoded in 7-qubit Steane codes connected via a transversal cnot. The uncorrectable error rate has probability $0.088\pm 0.004$, compared to total error rate $0.335\pm 0.003$ (to 2 standard deviations).
@@
With this joint probability distribution, we can compute logical error rates associated with the transversal cnot by classifying errors into correctable and uncorrectable, in accordance with the properties of the Steane code. The Steane code, having distance $d=3$, can correct one arbitrary weight-1 error on a single logical code block. Furthermore, since it is a CSS code it may correct weight-2 errors comprised of $X$ and $Z$, namely $X_{i}Z_{j}$ for $i\neq j$. Here, we look in particular at two copies of the 7-qubit Steane code that are connected via a transversal cnot. In that context, errors remain correctable so long as they satisfy these conditions independently on either code block. This means all errors supported on a single cnot are correctable, since it remains weight-1 on each code block under the action of the cnot. Additionally, any weight-2 $X_{i}Z_{j}$ over both code blocks is correctable, even though it can spread via the transversal cnot to a higher weight term, because the error remains at most a single $X$ and single $Z$ on each respective code block. All errors that are stabilizer-equivalent to these are correctable; all other errors are uncorrectable. Whenever an orbit contains mixed-weight terms, if any term in the orbit is uncorrectable then we classify the error as uncorrectable, to give the worst-case error.

### v2

# Also at: ]Alpine Quantum Technologies GmbH, Innsbruck, Austria

@@ II.1 Setting and conventions
Figure 1: Cycle benchmarking circuit design [14] leveraged to learn individual Pauli eigenvalues to relative precision. Illustrated in this example is the simpler form that is sufficient for any Clifford hard cycle $\mathcal{H}$ where easy cycles are Pauli randomized cycles $P$ [15]. The easy cycle $B$ includes local rotations to prepare Pauli product states and define the measurement basis. The randomized Pauli cycle AA eliminates bias from measurement errors.
@@
We will also consider the restriction of elements in $\mathbb{P}_{n}$ to arbitrary subsets of qubits. The restriction of Pauli $P$ to the qubits indexed by elements of set $S$ is denoted $P_{S}$. For example, $Z_{1}X_{2}Y_{4}]_{\{1\}}=Z_{1}$, $Z_{1}X_{2}Y_{4}]_{\{2,3\}}=X_{2}I_{3}$, or $Z_{1}X_{2}Y_{4}]_{\{4,1\}}=Y_{4}Z_{1}$. Note that the ordering of indices in the set is used here to typographically order the restricted Pauli. Similarly, we denote by $\mathbb{P}_{S}$ the subgroup of $\mathbb{P}_{n}$ that is supported on only the qubits of $S$. It will also be useful in this context, to define the map $\omega$ between arbitrary Pauli operators $P$ and $Q$, defined as
@@
where $p_{i}=\chi_{ii}$ is the classical probability of some Pauli operator $P_{i}$ occurring in the quantum channel. Evidently, the number of non-trivial parameters in $\mathcal{P}$ is drastically smaller than in $\mathcal{G}$; at most $4^{n}-1$. These smaller operators are often more practical to study.
@@
Channels of interest are generally hard cycles that are not the identity or idling operation. Typically we have some ideal channel $\mathcal{H}$ that is implemented imperfectly as $\tilde{\mathcal{H}}$$\mathcal{E}\circ\mathcal{H}$ where $\mathcal{E}$ is the noise channel we are interested in quantifying. We are thus not always able to insert random Pauli operators either side of the noise channel $\mathcal{E}$, but we may still interleave operations either side of $\tilde{\mathcal{H}}$. We can effectively twirl $\mathcal{E}$ if we instead insert a Pauli operator $P_{i}^{c}=\mathcal{H}^{\dagger}(P_{i})$ before $\tilde{\mathcal{H}}$ for each random $P_{i}$ inserted after $\tilde{\mathcal{H}}$.
@@ II.3 Probability Marginalization
$\mu_{S}(P)=\sum_{\begin{subarray}{c}P_{i}\in\mathbb{P}_{n}\\ [P_{i}]_{S}=P\end{subarray}}p(P_{i}),$ (7)
where we remind the reader that $P_{i}$ is an $n$-qubit Pauli operator in $\mathbb{P}_{n}$, and $P_{i}]_{S}=P$ denotes that the restriction of $P_{i}$ to the subset $S$ must equal $P$. Generically we refer to a marginal probability distribution over $n$ qubits as ‘$n$-qubit’ marginals.
@@
By collating marginal probabilities and unique eigenvalues into vectors, $\bm{\mu}$ and $\bm{\lambda}$, we can straightforwardly express equations like Eqs. (4) and (9) in the form $\bm{\mu}=W\bm{\lambda}$. If $\bm{\lambda}$ consists of all Pauli eigenvalues in $\mathbb{P}_{n}$$W$ is the Walsh-Hadamard Transform. More generally, Eq. (9) allows for $W$ to have elements that are zero, and may cause certain eigenvalues to contribute asymmetrically to different marginal probabilities.
@@ II.4 Graphical model under Gibbs random field
The error correlations of single-qubit and entangling gates realized in the used trapped-ion experiment are assumed to be nearest-neighbor. We can abstract to a Gibbs random field connectivity graph as discussed in Sec. IV, which we will use to make predictions about logical performance. Each qubit is connected to its partner in the entangling gate and the corresponding neighbors.
The considerations above are tailored to our experiment and may not hold in general. Other platforms may be dominated by different error sources and error correlations, and the details of the relevant GRF model and its compactness will vary. Regardless, this GRF framework can be applied generally.
@@ II.5 Optimizing Experimental Parameters
In the same vein, we want to learn chosen experimental parameters in as few experimental runs as possible. One can extract up to $2^{n}$ different Pauli expectation values in parallel, but generally this requires entangled input states and measurements which contribute to SPAM. However, for hard cycles from the Clifford group certain finite sequence lengths are equivalent to performing the identity operation. Our hard cycle is a transversal application of cnots, so we specifically have $\mathcal{H}^{2}(\rho)=\rho$. Hence, by choosing $\mathcal{H}^{m}$ such that $m$ is even, we guarantee that easily-prepared Pauli product states will be mapped to Pauli product measurements, keeping the contribution of SPAM low. At the same time, the measurements resolve each qubit and we can infer multi-qubit Pauli measurement outcomes by taking the product of single-qubit Pauli measurement outcomes. This reduces the number of experiments further because we are able to extract all $2^{n}$ distinct Pauli expectation values from $n$-qubit product states.
Figure 3: Optimization of experimental resources. a The number of distinct experiments to run for a single, two, and seven parallel cnots to obtain all unique orbits is 4 (the eigenstates of elements from the set AA, for 1-cnot orbits), $4\times 9=36$ (all combinations of $A$ and $B$, for 2-cnot orbits), and $36\times 3=108$ (all tilings $C$ of elements from AA and $B$, for all possible 2-cnot orbits over 7 cnots), respectively. b Tradeoff between number of randomizations per initial state and shots per randomization for an eigenvalue estimate of $\lambda_{ZX}$. The color scale depicts the standard deviation in estimates of $\lambda_{ZX}$ at the chosen experimental parameters, estimated by subsampling over an experimental data set. In total the data set included 300 randomizations with 650 shots each. In our platform changing the randomization is more costly than repeating a shot. Time in minutes to estimate a fixed initial state using three values for $m$ is indicated by red contours. Total shots required is obtainable by the product of the shots per randomization, total randomizations, the distinct sequence lengths $m$, and the distinct initial states.
@@
A naive estimate for the total number of configurations needed for 2-cnot marginals would be $4\times 4=16$ initial states. However, the set $(P_{1}P_{2})^{\circlearrowright\mathcal{H}}\neq P_{1}^{\circlearrowright\mathcal{H}}\times P_{2}^{\circlearrowright\mathcal{H}}$. This can be observed simply by noting the cardinality of the sets: $(P_{1}P_{2})^{\circlearrowright\mathcal{H}}=\{P_{1}P_{2},\mathcal{H}(P_{1}P_{2})\}$ has an orbital length of two but $P_{1}^{\circlearrowright\mathcal{H}}\times P_{2}^{\circlearrowright\mathcal{H}}$ is a set of 4 elements. For our choice of hard cycle, equality is satisfied whenever $P_{1}$ or $P_{2}$ is mapped to itself, up to a change in sign. Consequently we often need to prepare $P_{1}$ and $P_{1}P_{2}$, increasing the number of initial states from 16 to 36. When considering all possible 2-cnot marginalizations in the transversal seven cnot case, three distinct tilings, shown in Fig. 3 a, are needed such that 100 initial states are required in total ($3\times 36=108$, minus 8 redundancies) to learn $4^{4}\times\binom{7}{2}=5376$ expectation values.
@@ II.6 Uncertainties and noise floors
Error bars shown in figures depicting marginal error rates are determined by bootstrapping the data over the time period of the experiment. Where calculated, the range is reported to one standard error unless stated otherwise. Such bars do not bound the long-term performance of the device, rather, they signify the variance in the experimental data over the time-frame that the data was obtained.
@@ III Experimental setup
The Coulomb interaction between the ions gives rise to collective motional modes of the ions, which are used to mediate entangling operations between any desired pair of qubits. Single-ion addressing in the radial chain direction with light at the $729\text{\,}\mathrm{nm}$ qubit transition allows for arbitrary single-qubit and arbitrary-pair two-qubit gate operations. The harmonic potential confining the ions leads to inter-ion spacings that are non-uniform such that addressing errors like optical crosstalk depend on which ions are addressed, as seen in Fig. 4 a. More details on the setup can be found in Ref. [24].
The native gate set used in the setup consists of arbitrary-angle single-qubit rotations $R(\theta,\phi)$ around any axis in the equatorial plane of the Bloch sphere (pulse area $\theta$ and angle $\phi$ relative to +X+X) implemented via laser pulses resonant with the optical qubit transition, single-qubits $Z$-rotations ‘virtually’ implemented in software, and maximally-entangling two-qubit gates $MS(\pm\pi/2)=\mathrm{exp}(\mp i\frac{\pi}{4}X\otimes X)$ implemented via the Mølmer-Sørensen interaction [41]. The $MS$ gate belongs to the class of geometric phase gates generated by optical forces where entanglement between qubits not natively interacting is created via a mediating boson; here excitations of the joint motional mode of the ion crystal. The $MS(\pm\pi/2)$ gate is equivalent to the cnot gate up to local rotations [42], compiled here as
@@
The $MS$ gates employed here utilize the center-of-mass radial motional mode of the ion crystal as the mediating boson. Using the global phononic bus of the motion to create entanglement means that we have to perform $MS$ gates serially even when two gates feature disjoint support. There are control-theoretic gate modulation techniques to circumvent this limitation, e.g. in Ref. [43], which are however not part of this work.
@@ IV.2 Transversal CNOT
Figure 6: Characterizations of a 14-qubit transversal cnot $\mathcal{H}=\prod_{i=0}^{6}\textsc{cnot}_{i,i+9}$ in a 16-ion register. a Heatmap display of 1-cnot marginals for each constituent cnot. The $ZX$ error terms are identified to be caused by a miscalibration of the $MS$ gate power. b A heatmap display of 1-cnot marginals for a separate characterization experiment, after the source of the ZXZX signal was identified and eliminated. In a and b, marginals $\mu_{S}(ZI^{\circlearrowright\mathcal{H}})$ and $\mu_{S}(IZ^{\circlearrowright\mathcal{H}})$ associated with dephasing are larger than for a single cnot in isolation. The error characteristics for each of the individual cnots is similar (little change horizontally) with the exception of cnot0,9 for $\mu_{S}(IY^{\circlearrowright\mathcal{H}})$ and $\mu_{S}(IZ^{\circlearrowright\mathcal{H}})$ (highlighted in red). c Significant error terms for a 2-cnot marginal, taken at the same time as b, over cnot5,14 and cnot6,15. Marginal terms only supported on $(6,15)$ are shown first, then terms only supported on $(5,14)$, and then all other terms with error rates higher than the dark-gray region. d The same seven orbits from b as histograms, with the anomalous strengths of cnot0,9 again highlighted in red. Apart from the anomalous terms in cnot0,9, there is a much higher amount of dephasing on control and target qubits due to time spent idling. Since the predominant difference from single cnots in isolation appears to be idling, we compare the noise profile of individual cnots from the transversal cnot of d with a single cnot undergoing an equivalent period of idling. e Experimental sequence for labels ‘delay’ and ‘seven’ in f. The top circuit executes seven cnots successively and we study the marginal on support $(0,9)$ and $(3,12)$. The bottom circuit executes a single cnot but idles for an additional time equivalent to 6 cnots being applied in sequence. f and g are histograms of the marginals over cnot0,9 and cnot3,12, respectively, either as part of the full transversal cnot or as a single cnot with an equivalent idling time of the transversal cnot. Data in g was taken with machine performance deteriorated before ongoing larger maintenance. The idling errors of f resemble the transversal cnot errors, rather than additional dephasing.
@@
The overall wall-clock runtime for CER of a transversal cnot suite with 20 randomizations per initial state and 200 shots per randomization on seven supports takes about 40 hours after optimization; this time includes automatic calibration routines and other overhead of operation.
@@ IV.3 Predicting logical error rates and performance
Figure 7: Probabilities of different error scenarios reconstructed from transversal cnot CER data using the GRF factor graph of Fig. 4 c. Reconstructed non-zero error rates are grouped into uncorrectable and correctable errors (including errors stabilizer-equivalent to correctable, see Sec. II.4) assuming two logical qubits encoded in 7-qubit Steane codes connected via a transversal cnot. The uncorrectable error rate has probability $0.088\pm 0.004$, compared to total error rate $0.335\pm 0.003$. Error bars are obtained by bootstrapping the experimental data to generate marginal error rates, as outlined in Sec. II.6. The reconstructed rates are reported here to two standard errors.
@@
With this joint probability distribution, we can compute logical error rates associated with the transversal cnot by classifying errors into correctable and uncorrectable, in accordance with the properties of the Steane code. The Steane code, having distance $d=3$, can correct one arbitrary weight-1 error on a single logical code block. Furthermore, since it is a CSS code it may correct weight-2 errors comprised of $X$ and $Z$, namely $X_{i}Z_{j}$ for $i\neq j$. Here, we look in particular at two copies of the 7-qubit Steane code that are connected via a transversal cnot. In that context, errors remain correctable so long as they satisfy these conditions independently on either code block. This means all errors supported on a single cnot are correctable, since it remains weight-1 on each code block under the action of the cnot. Additionally, any weight-2 $X_{i}Z_{j}$ over both code blocks is correctable, even though it can spread via the transversal cnot to a higher weight term, because the error remains at most a single $X$ and single $Z$ on each respective code block. All errors that are stabilizer-equivalent to these are correctable; all other errors are uncorrectable. Whenever an orbit contains mixed-weight terms, if any term in the orbit is uncorrectable then we classify the error as uncorrectable, to give the worst-case error. The ability to distinguish fine-grained errors, such as an uncorrectable $X_{0}X_{1}$ from a correctable $X_{0}X_{9}$, is available to us through CER and is what ultimately allows us to obtain logical estimates with multiplicative precision.